# An alternative theory of magnetic flux tubes in strong fields via axion origin photons

### v2

@@ I Introduction
Figure 8: Topological effects of magnetic reconnection in the lower (left) or upper (right) layers of the magnetic tube. Here, the unipolar part of the $\Omega$-loop is rebuilt on its base, compressing the $\Omega$-loop (blue lines) to form a free O-loop (green lines) (Fig. 5a). The yellow lines show the movement of the substance leading to the connection of the “leg” loop (see analogous Fig. 4 in E. N. Parker (1994)). If the O-loop (primary magnetic reconnection, green lines) can randomly have different directions of magnetic fields, then the secondary magnetic reconnection can generate loop “legs” in different layers, for example, in lower layers (Fig. 8c and also Fig. 8d as an analog of Fig. 5b) and upper layers (Fig. 8e and also Fig. 8f as an analog of Fig. 4 in S. Jabbari et al. (2016)).
From here, since the sunspot cycle is identical to the maxima and minima of the magnetic tube number on the surface of the Sun, we are interested in the physics of the maxima and minima both during the primary reconnection for the time $(\tau_{d})_{conv}$ of the rise of the almost empty magnetic tube from the tachocline to the solar photosphere, which are associated with the features of the tilt angle of Joy’s law (see Fig. 10c), and during the secondary reconnection for the time $(\tau_{n})_{rec}$ of the departure of the same rise of the almost empty magnetic tube from the surface of the Sun and, as a consequence, their “disappearance” on the Sun.
II.2.1 On the ascending magnetic tube and the formation of sunspots both near the tacho-cline with $\sim 10^{7}~G$ in the absence of Joy’s law and in the middle region with $\sim 10^{5}~G$ in the presence of Joy’s law
@@
For the neutral buoyancy condition ($\rho_{int}=\rho_{ext}\approx 0.2~g/cm^{3}$; see Figs. 3 and 6) and the strong toroidal field of the magnetic tube, $B_{Sun}=4.1\cdot 10^{7}~G$, as well as the average width of the “thin” ring $a_{conv}\sim 3.7\cdot 10^{-4}~H_{p}$ (see Eq. (22) and Fig. 6) between the O-loop and the walls of the magnetic tube (see Fig. 6), it is easy to show the estimate of the time $(\tau_{d})_{conv}$ (see Eqs. (27)-(34)) and the velocity $(v_{rise})_{conv}$ (see Eq. (16) and also Eqs. (36)-(37) of the convective rise from the tachocline to the photosphere, which for such large magnetic fields has a significant amount of rise of the magnetic tubes at all latitudes (see Fig. 5a and the red lines in Fig. 9a).
@@
Figure 9: Scheme of turbulent reconstruction of the toroidal magnetic field in the convective zone (CZ): (a) meridional circulation (closed blue poloidal field lines), which is the source of the toroidal field in the tachocline (thin black lines) based on the holographic mechanism of quantum gravity (see Fig. 1b), and the magnetic buoyancy of the flux tubes (red arrows) with a magnetic field of $\sim 10^{7}~G$; (b) In the CZ, through secondary turbulent reconnection of the magnetic tube near the tachocline (see Fig. 10b2,c), effects of both macroscopic turbulent plasma diamagnetism are observed, facilitating the formation of two layers on the surface (see analogous (V. N. Krivodubskij, 2005, 2021; V. N. Krivodubskii, 1992)) with different vertical directions of horizontal field transfer (see “black” arrows), and the conditions of magnetic field transfer along $\nabla\rho$ due to small-scale magnetic pulsations (the so-called magnetic pumping, see “white” arrows), acting against buoyancy. However, in the deep layers in the near-equatorial region, the $\nabla\rho$ effect provides upward advection, which, on the contrary, directs magnetic buoyancy, and thus, can facilitate the penetration of strong fields to the surface.
@@
This solution clearly depends on the rise time of the magnetic tubes rising from the tachocline to the solar surface. Therefore, the primary magnetic reconnection itself in the lower layers of the magnetic tubes (see Fig. 5a), is not the final stage of the modeling. The essence of a practically empty tube in strong fields of $\sim 10^{7}~G$ is related (through the primary magnetic reconnection) to the physics of the secondary turbulent reconnection of magnetic bipolar structures (see Fig. 5b, and also Fig. 4 in (A. S. Brun et al., 2011)). Here, on its basis, the bipolar part of the $\Omega$-loop is reconstructed, compressing the $\Omega$-loop (blue lines: see Fig. 5b,c) near the tachocline and simultaneously lifting them (from 2 to 3 years) to the surface of the Sun, and, as a consequence, after some time (from several days to several months (see S. K. Solanki (2003); K. Petrovay & L. van Driel-Gesztelyi (1997)) a free O-loop appears (blue lines in Fig. 5d), which eventually “disintegrates” or, more precisely, the O-loop (without spots) flies away from the surface of the Sun (see Fig. 5d).
@@
In this regard, we obtain an answer to this important question. For this purpose, we consider the second part of the scenario, based on the remarkable idea of L. L. Kichatinov (1991), which includes the generation of the magnetic field of $\leqslant 2\cdot 10^{5}~G$ near the bottom of the convective zone (see Fig. C.1b in V. D. Rusov et al. (2021) for the holographic mechanism) and transfer of the toroidal field from the deep layers at different latitudes.
It is very important that the efficiency of the magnetic buoyancy transfer is associated with the participation of two processes: macroscopic turbulent diamagnetism (Y. B. Zel’dovich, 1957, 1956; L. J. Spitzer, 1957) and rotational $\nabla\rho$-pumping (L. L. Kichatinov, 1991).
Let us note some important properties of macroscopic turbulent diamagnetism. It is known that Y. B. Zel’dovich (1956, 1957) and L. J. Spitzer (1956) discovered the diamagnetism of inhomogeneously turbulent conducting liquids, in which the inhomogeneous magnetic field moves as a single whole. In this case the turbulent fluid, for example, with nonuniform effective diffusivity $\eta_{T}\approx(1/3)vl$ (see Fig. 1 in  L. L. Kitchatinov & G. Rüdiger (2008), and also (Y. B. Zel’dovich, 1957, 1956; L. J. Spitzer, 1957)) behaves like a diamagnetic one and carries the magnetic field with the effective velocity (H. Köhler, 1973)
@@
where $l$ is the mixing length of turbulent pulsations, and $v=\sqrt{\langle v^{2}\rangle}$ is the root-mean-square velocity of turbulent motion. The minus sign on the right in Eq. (39) shows the meaning of turbulent magnetism: it is not paramagnetic magnetism, so magnetic fields repel from regions with relatively high turbulent intensity. In other words, macroscopic turbulent plasma diamagnetism and, as a consequence, the so-called macroscopic diamagnetic effect (see K.-H. Rädler (1968a, b)) in the physical sense is the displacement of the averaged magnetic field $B$ from regions with increased intensity of turbulent pulsations to regions with less developed turbulence (F. Krause & K.-H. Raedler, 1980; L. L. Kichatinov & G. Ruediger, 1992).
@@
However, there is an interesting problem of the diamagnetic process caused by inhomogeneous turbulent intensity with allowance for the total nonlinearities in the magnetic field. This is due to the fact that up to the present time analytical estimates have been obtained only for the limiting cases of weak and strong magnetic fields. For example, at strong magnetic fields of flux tubes, at $B>10^{5}~G$, e.g. $\geqslant 10^{7}~G$, the diamagnetic effect becomes almost negligible, in particular, strong magnetic damping of diamagnetism $\sim B^{-3}$ is obtained for super-equipartitions of fields (F. Chen et al., 2022) when turbulence is close to two-dimensional (Y. B. Zel’dovich, 1957). On the other hand, for very weak fields the diamagnetic pumping, which is predetermined by the intensity of turbulence at $B<10^{5}~G$, is a very effective process (F. Chen et al., 2022).
@@
Among the known limiting cases of weak and strong magnetic fields from the tachocline to the surface, we are interested in the addition to the portions of strong fields (via the secondary Lazaryan-Vishniac reconnection (see A. Lazarian & E. T. Vishniac (1999); A. Lazarian et al. (2004); G. Kowal et al. (2009), and also §III and §IX in (A. Lazarian et al., 2020)) near the tachocline, which have a very slow Sweet-Parker velocity $(V_{rec})_{2}$ (see E. N. Parker (1957); P. A. Sweet (1958)) for Fig. 10b2), the part of the weak magnetic field from $0.8~R_{Sun}$ to near $0.85~R_{Sun}$ (see Fig. 10b2,c), which constitutes the super-equipartitions of fields $B\geqslant B_{eq}$ and is the source of two processes: diamagnetic pumping, and rotational $\nabla\rho$-pumping. This is due to the fact that the secondary magnetic reconnection of the flux tubes (see Figs. 8b,d,f,g) leads to the real increase in the upward magnetic field to $B\approx 2\cdot 10^{5}~G$ at $\sim 0.8~R_{Sun}$, which is identical to the $1/\beta=B^{2}/8\pi p_{ext}\approx 2.9\cdot 10^{-5}$ (see e.g. Eq. (9), and Fig. 10b2,c). This means that the real toroidal magnetic flux tube is a consequence of the formation of the secondary reconnection, which generates both the “anti-buoyant” effects of downwardly directed turbulent diamagnetic transport and the rotational effect of the magnetic $\nabla\rho$-pumping at polar latitudes, and the “buoyant” effects of magnetic flux tubes at equatorial latitudes as a result of the upward rotational effect of magnetic $\nabla\rho$-pumping.
@@
Figure 10: An alternative theory of magnetic flux tubes in strong fields by means of axion origin photons associated with primary and secondary reconnections.
@@
Figure 11: (cont.) An alternative theory of magnetic flux tubes in strong fields by means of axion origin photons associated with primary and secondary reconnections.
@@
In this regard, we consider turbulence with quasi-isotropic spectral tensor (L. L. Kichatinov, 1987, 1991), which is certainly the simplest representation for inhomogeneous turbulence (see Eq. (2.12) in L. L. Kichatinov & G. Ruediger (1992)). As a consequence, the information on the spectral properties of turbulence (given by Eqs. (2.12)-(2.16) from L. L. Kichatinov & G. Ruediger (1992)) is sufficient to reduce the expression for the average electromotive force $\varepsilon$ (see Eq. (2.1) L. L. Kichatinov & G. Ruediger (1992)) to its traditional form, where only integrations over the wave number $k$ and frequency $\omega$ remain. After such shortening it is possible to get (see Eq. 3.1 in L. L. Kichatinov & G. Ruediger (1992))
@@
We are interested in the problem of reconstructing a strong toroidal field of flux tubes $\sim 10^{7}~G$ (see Figs. 9a,b with the magnetic tube profile, and Figs. 10a,b,c with the magnetic tube in full face), transform regions of the average magnetic field $B\sim 2\cdot 10^{5}~G$ in the convective zone (see Figs. 10b,c), and thereby allow the organization of the amazing balance between the magnetic buoyancy, turbulent diamagnetism, and the rotationally modified $\nabla\rho$-effect.
@@
Here we assumed that, according to Figs. 10b2,c, the radial profile of turbulent diffusion is a smooth-convex function with a maximum, $\eta_{T}\approx 10^{11}~cm/s$, approximately in the middle of the CZ at a depth of $z=1.30\cdot 10^{5}~km$ at $\sim 0.8R_{Sun}$, which is described by the distance of $2.3\cdot 10^{4}~km$ between $0.8R_{Sun}$ and near $0.85R_{Sun}$ (see Figs. 10b2,c) and with the depth of the bottom of the convective zone $z_{0}\approx 1.84\cdot 10^{5}~km$ (see M. Stix (2002); R. Kippenhahn et al. (2012); J. Schumacher & K. R. Sreenivasan (2020)).
In order to consider the balance of the three velocities $v_{B}$, $v_{dia}$, and $v_{dens}$, it is necessary to apply the widely used approximation of the mixing length (see e.g. E. Böhm-Vitense (1958); P. Bradshaw (1974); D. O. Gough (1977); D. Gough (1977); A. J. Barker et al. (2014); A. Brandenburg (2016)), which, according to L. L. Kichatinov (1991), fully satisfies this goal. This approximation will be understood as the replacement of nonlinear terms along with time derivatives in the equations for fluctuating fields by means of $\tau$-relaxation terms, i.e. instead of equations (3.1) and (3.9) from L. L. Kichatinov (1991), we now have the equation of the radial speed of the toroidal field in the convection zone
@@
where $\tau\approx l/(\langle u^{2}\rangle^{\circ})^{1/2}$ is a typical lifetime of a convective eddy; $l$ is the mixing length; $\langle u^{2}\rangle^{\circ}$ is the mean intensity of fluctuating velocities for original turbulence; $\theta$ is the latitude (scalar); $e_{r}\nabla\rho/\rho=-e_{r}g/[(\gamma-1)c_{p}T]$ , where $e_{r}$ is the radial unit vector; $T\cong 1.352\cdot 10^{6}~K$ is the temperature at $0.8R_{Sun}$ (J. Bahcall & M. Pinsonneault, 1992); $g=g_{0}(R_{Sun}/r)$ is the gravity, where $g_{0}=2.74\cdot 10^{4}~cm/s^{2}$ is the surface value of solar gravity, and $g_{0.8R_{Sun}}=4.22\cdot 10^{4}cm/s^{2}$ at $0.8R_{Sun}$; $c_{p}=3.4\cdot 10^{8}~cm^{2}s^{-2}K^{-1}$ (fully ionized hydrogen) is the specific heat at constant pressure; $\gamma=5/3$ is the ratio of specific heats $c_{p}/c_{V}$; $3\kappa=\tau\langle u^{2}\rangle^{\circ}$, where $\kappa\equiv\eta_{T}\approx 10^{11}~cm^{2}/s$ is turbulent diffusivity coefficient, and the mixing length relation $\langle u^{2}\rangle^{\circ}=-\nabla\Delta Tl^{2}g/(4T)$, where the $\nabla\Delta T$ is superadiabatic temperature gradient; $v_{p}=(1/6)\tau\langle u^{2}\rangle^{\circ}(\nabla\rho/\rho)$ is the velocity of the magnetic field transfer caused by the density gradient (see Eq. (36) in S. I. Vainshtein & L. L. Kichatinov (1983)); $\hat{\Omega}=Co=2\tau\Omega$ is the Coriolis number (reciprocal of the Rossby number), where $\Omega$ is the rotation speed, $\tau$ is the turnover time, and the functions
@@
(see also analogous Eq. (19) and Fig. 2 in L. L. Kitchatinov & A. A. Nepomnyashchikh (2016)), which describe the rotational effect on turbulent convection.
@@
Thirdly, in addition to the radial profile of turbulent diffusion $\eta_{T}\approx 10^{11}~cm^{2}/s$, which is related to the diamagnetic pumping velocity $v_{dia}\approx 1.7\cdot 10^{2}~cm/s$ and the distance $2.3\cdot 10^{4}~km$ between $0.8R_{Sun}$ and near $0.85R_{Sun}$ (see the “green” line of the turbulent-toroidal magnetic field in Fig. 10b,c), we need to calculate the magnetic buoyancy velocity $v_{B}$, which is associated with a super-uniformly distributed field strength $B\sim 2\cdot 10^{5}~G$ (see Fig. 9b and Fig. 10b,c). It should be remembered that for buoyant flux tubes with super-equipartition field strength $B\sim 2\cdot 10^{5}~G\geqslant(H_{p}/a)^{1/2}B_{eq}\approx 3.16B_{eq}$ with $a\approx 0.1H_{p}$, where, as Yuhong Fan wrote (see Section 7 in (Y. Fan, 2021)), $B_{eq}\approx 6.1\cdot 10^{4}~G$ is in equipartition with the kinetic energy density of the convective downflows, the magnetic buoyancy of the tubes dominates the hydrodynamic force from the convective downflows and the flux tubes can rise unaffected by convection”, but significantly affects the turbulent diamagnetic pumping!

### v3

@@ II Physics of the theory of practically empty flux tubes in strong fields of the order $\sim 10^{7}~G$
From here, since the sunspot cycle is identical to the maxima and minima of the magnetic tube number on the surface of the Sun, we are interested in the physics of the maxima and minima both during the primary reconnection for the time $(\tau_{d})_{conv}$ of the rise of the almost empty magnetic tube from the tachocline to the solar photosphere (see about $\sim 1-2$ days at max in Eq. (31) and about $\sim 10$ days at min in Eq. 35), and during the secondary reconnection $(\tau_{n})_{rec}$, which are associated with the features of the tilt angle of Joy’s law (see Fig. 8c), during the rise of the almost empty magnetic tube from the surface of the Sun (see about $\sim 2$ years in Fig. 8b2 and c) and, as a consequence, their “disappearance” on the Sun (see about <0.5<0.5 years in Fig. 8c).
II.2.1 Secondary reconnection of the MFT without axions, but with concern of turbulent diffusion and the tilt angle of the Joy law
@@
For the condition of neutral buoyancy ($\rho_{int}=\rho_{ext}\approx 0.2~g/cm^{3}$ [49]; see A.16 in [14], Figs. 2 and 5) and the strong toroidal field of the magnetic tube, $B_{Sun}=4.1\cdot 10^{7}~G$, as well as the condition of the average width of the “thin” ring $a_{conv}^{min}\sim 3.7\cdot 10^{-4}~H_{p}$ (see Eq. (24)) between the O-shaped loop and the walls of the magnetic tube (see Fig. 5), it is easy to recall the estimate of the time $(\tau_{d})_{conv}^{max}\sim 1.3~days$ (see Eq. (31)) and velocity $(v_{rise})_{conv}^{max}\sim 1.40~km/s$ (see Eq. (33), as well as the time $(\tau_{d})_{conv}^{min}\sim 10~days$ (see Eq. (35)) and velocity $(v_{rise})_{conv}^{min}\sim 0.14~km/s$ (see Eq. (34) of the convective rise from the tachocline to the photosphere, which for such large magnetic fields has a significant amount of rise of the magnetic tubes at all latitudes (see Fig. 4a and the red arrows of the magnetic flux tubes in Fig. 7a).
@@
Figure 7: Scheme of turbulent reconstruction of the toroidal magnetic field in the convective zone (CZ): (a) meridional circulation (closed blue poloidal field lines), which is the source of the toroidal field in the tachocline (thin black lines), and the magnetic buoyancy of the flux tubes (red arrows) with a magnetic field of $\sim 10^{7}~G$; (b) In the CZ, through secondary turbulent reconnection of the magnetic tube near the tachocline (see Fig. 8b2,c), effects of both macroscopic turbulent plasma diamagnetism are observed, facilitating the formation of two layers on the surface (see analogous [80, 81, 82]) with different vertical directions of horizontal field transfer (see “black” arrows), and the conditions of magnetic field transfer along $\nabla\rho$ due to small-scale magnetic pulsations, acting against buoyancy. However, in the deep layers in the near-equatorial region, the $\nabla\rho$ effect provides upward advection, which, on the contrary, directs magnetic buoyancy, and thus, can facilitate the penetration of strong fields to the surface.
@@
This solution clearly depends on the rise time of the magnetic tubes rising from the tachocline to the solar surface. Therefore, the primary magnetic reconnection itself in the lower layers of the magnetic tubes (see Fig. 4a), is not the final stage of the modeling. The essence of a practically empty tube in strong fields of $\sim 10^{7}~G$ is related (through the primary magnetic reconnection) to the physics of the secondary turbulent reconnection of magnetic bipolar structures (see Fig. 4b, and also Fig. 4 in [60]). Here, on its basis, the bipolar part of the $\Omega$-loop is reconstructed, compressing the $\Omega$-loop (blue lines: see Fig. 4b,c) near the tachocline and simultaneously lifting them (from 2 to 3 years) to the surface of the Sun, and, as a consequence, after some time (from several days to several months (see [83, 84]) a free O-loop appears (blue lines in Fig. 4d), which eventually “disintegrates” or, more precisely, the O-loop (without spots) flies away from the surface of the Sun (see Fig. 4d).
@@
In this regard, we obtain an answer to this important question. For this purpose, we consider the second part of the scenario, based on the remarkable idea of L.L. Kitchatinov [84], which includes the generation of the magnetic field of $\leqslant 2\cdot 10^{5}~G$ near the bottom of the convective zone and transfer of the toroidal field from the deep layers at different latitudes. It is very important that the efficiency of the magnetic buoyancy transfer is associated with the participation of two processes: macroscopic turbulent diamagnetism [85, 86, 87] and rotational $\nabla\rho$-pumping [84].
Let us note some important properties of macroscopic turbulent diamagnetism. It is known that Zel’dovich [86, 85] and Spitzer [87] in 1957 discovered the diamagnetism of inhomogeneously turbulent conducting liquids, in which the inhomogeneous magnetic field moves as a single whole. In this case the turbulent fluid, for example, with nonuniform effective diffusivity $\eta_{T}\approx(1/3)vl$ (see Fig. 1 in [88], and also [85, 86, 87]) behaves like a diamagnetic one and carries the magnetic field with the effective velocity [89]
@@
where $l$ is the mixing length of turbulent pulsations, and $v=\sqrt{\langle v^{2}\rangle}$ is the root-mean-square velocity of turbulent motion. The minus sign on the right in Eq. (36) shows the meaning of turbulent magnetism: it is not paramagnetic magnetism, so magnetic fields repel from regions with relatively high turbulent intensity. In other words, macroscopic turbulent plasma diamagnetism and, as a consequence, the so-called macroscopic diamagnetic effect (see [90, 91]) in the physical sense is the displacement of the averaged magnetic field $B$ from regions with increased intensity of turbulent pulsations to regions with less developed turbulence [92, 93].
@@
However, there is an interesting problem of the diamagnetic process caused by inhomogeneous turbulent intensity with allowance for the total nonlinearities in the magnetic field. This is due to the fact that up to the present time analytical estimates have been obtained only for the limiting cases of weak and strong magnetic fields. For example, at strong magnetic fields of flux tubes, at $B>10^{5}~G$, e.g. $\geqslant 10^{7}~G$, the diamagnetic effect becomes almost negligible, in particular, strong magnetic damping of diamagnetism $\sim B^{-3}$ is obtained for super-equipartitions of fields [94] when turbulence is close to two-dimensional [85]. On the other hand, for very weak fields the diamagnetic pumping, which is predetermined by the intensity of turbulence at $B<10^{5}~G$, is a very effective process [94].
@@
Among the known limiting cases of weak and strong magnetic fields from the tachocline to the surface, we are interested, in the addition to the portions of strong fields (via the secondary Lazaryan-Vishniac reconnection (see  [95, 96, 97], and also §III and §IX in [98]) near the tachocline, which have a very slow Sweet-Parker velocity $(V_{rec})_{2}$ (see [99, 100]) for Fig. 8b2), the part of the weak magnetic field from $0.8~R_{Sun}$ to near $0.85~R_{Sun}$ (see Fig. 8b2,c), which constitutes the super-equipartitions of field $B\geqslant B_{eq}$ and is the source of two processes: diamagnetic pumping, and rotational $\nabla\rho$-pumping. This is due to the fact that the secondary magnetic reconnection of the flux tubes (see Figs. 6b,d,f,g) leads to the real increase in the upward magnetic field to $B\approx 2\cdot 10^{5}~G$ at $\sim 0.8~R_{Sun}$, which is identical to the $1/\beta=B^{2}/8\pi p_{ext}\approx 2.9\cdot 10^{-5}$ (see e.g. Eq. (9), and Fig. 8b2,c). This means that the real toroidal magnetic flux tube is a consequence of the formation of the secondary reconnection, which generates both the “anti-buoyant” effects of downward turbulent diamagnetic transport and the rotational effect of the magnetic $\nabla\rho$-pumping at polar latitudes, and the “buoyant” effects of magnetic flux tubes at equatorial latitudes as a result of the upward rotational effect of magnetic $\nabla\rho$-pumping.
@@
Figure 8: An alternative theory of magnetic flux tubes in strong fields by means of axion origin photons associated with primary and secondary reconnections. [Ref: S.K. Solanki (2003) [101], K. Petrovay & L. van Driel-Gesztelyi (1997) [83].]
@@
Figure 8: (cont.) An alternative theory of magnetic flux tubes in strong fields by means of axion origin photons associated with primary and secondary reconnections. [Ref: A. Lazarian & E. T. Vishniac (1999) [95]; A. Lazarian et al. (2004) [96], (2020) [98]; G. Kowal et al. (2009) [97]; P. A. Sweet (1958) [100]; E. N. Parker (1955a) [6], (1957) [99], (1974) [11], (1974a) [12], (1974b) [10], (1977) [13]; V. D. Rusov et al. (2021) [14]]
@@
In this regard, we consider turbulence with quasi-isotropic spectral tensor [102, 84], which is certainly the simplest representation for inhomogeneous turbulence (see Eq. (2.12) in [93]). As a consequence, the information on the spectral properties of turbulence (given by Eqs. (2.12)-(2.16) from [93]) is sufficient to reduce the expression for the average electromotive force $\varepsilon$ (see Eq. (2.1) [93]) to its traditional form, where only integrations over the wave number $k$ and frequency $\omega$ remain. After such shortening it is possible to get (see Eq. 3.1 in [93])
@@
We are interested in the problem of reconstructing a strong toroidal field of flux tubes $\sim 10^{7}~G$ (see Figs. 7a,b for the magnetic tube profile, and Figs. 8a,b,c for the magnetic tube in full face), transform the regions of the average magnetic field $B\sim 2\cdot 10^{5}~G$ in the convective zone (see Figs. 8b,c), and thereby allow the organization of the amazing balance between the magnetic buoyancy, turbulent diamagnetism, and the rotationally modified $\nabla\rho$-effect.
@@
Here we assumed that, according to Figs. 8b2,c, the radial profile of turbulent diffusion is a smooth-convex function with a maximum, $\eta_{T}\approx 10^{11}~cm/s$, approximately in the middle of the CZ at a depth of $z=1.30\cdot 10^{5}~km$ at $\sim 0.8R_{Sun}$, which is described by the distance of $2.3\cdot 10^{4}~km$ between $0.8R_{Sun}$ and near $0.85R_{Sun}$ (see Figs. 8b2,c) and with the depth of the bottom of the convective zone $z_{0}\approx 1.84\cdot 10^{5}~km$ (see [103, 104, 105]).
In order to consider the balance of the three velocities $v_{B}$, $v_{dia}$, and $v_{dens}$, it is necessary to apply the widely used approximation of the mixing length (see e.g. [58, 106, 107, 108]), which, according to [84], fully satisfies this goal. This approximation will be understood as the replacement of nonlinear terms along with time derivatives in the equations for fluctuating fields by means of $\tau$-relaxation terms, i.e. instead of equations (3.1) and (3.9) from [84], we now have the equation of the radial speed of the toroidal field in the convection zone
@@
where $\tau\approx l/(\langle u^{2}\rangle^{\circ})^{1/2}$ is a typical lifetime of a convective eddy; $l$ is the mixing length; $\langle u^{2}\rangle^{\circ}$ is the mean intensity of fluctuating velocities for original turbulence; $\theta$ is the latitude (scalar); $e_{r}\nabla\rho/\rho=-e_{r}g/[(\gamma-1)c_{p}T]$ , where $e_{r}$ is the radial unit vector; $T\cong 1.352\cdot 10^{6}~K$ is the temperature at $0.8R_{Sun}$ [49]; $g=g_{0}(R_{Sun}/r)$ is the gravity, where $g_{0}=2.74\cdot 10^{4}~cm/s^{2}$ is the surface value of solar gravity, and $g_{0.8R_{Sun}}=4.22\cdot 10^{4}cm/s^{2}$ at $0.8R_{Sun}$; $c_{p}=3.4\cdot 10^{8}~cm^{2}s^{-2}K^{-1}$ (fully ionized hydrogen) is the specific heat at constant pressure; $\gamma=5/3$ is the ratio of specific heats $c_{p}/c_{V}$; $3\kappa=\tau\langle u^{2}\rangle^{\circ}$, where $\kappa\equiv\eta_{T}\approx 10^{11}~cm^{2}/s$ is turbulent diffusivity coefficient, and the mixing length relation $\langle u^{2}\rangle^{\circ}=-\nabla\Delta Tl^{2}g/(4T)$, where the $\nabla\Delta T$ is superadiabatic temperature gradient; $v_{p}=(1/6)\tau\langle u^{2}\rangle^{\circ}(\nabla\rho/\rho)$ is the velocity of the magnetic field transfer caused by the density gradient (see Eq. (36) in [109]); $\hat{\Omega}=Co=2\tau\Omega$ is the Coriolis number (reciprocal of the Rossby number), where $\Omega$ is the rotation speed, $\tau$ is the turnover time, and the functions
@@
(see also analogous Eq. (19) and Fig. 2 in [110]), which describe the rotational effect on turbulent convection.
From here, we obtain an the estimate of the Coriolis number for solar convection in the deep layer, which should be $\hat{\Omega}\approx 20$. By virtue of the equations (44)-(46) and the function $\phi_{n}$ (see Eq. (44)) for the value $\hat{\Omega}\approx 20$ (see from $0.8R_{Sun}$ in Fig. 8b2,c), the following values are adopted:
$\phi_{1}\cong 0.0171,~~~\phi_{2}\cong 0.0098,$ (47)
at which the radial velocity $v_{dens}^{light}$ in Eq. (40) of toroidal field transport changes the sign at the latitude $\theta^{*}=\arccos\sqrt{\varphi_{2}/\varphi_{1}}\cong 41^{\circ}$, being negative (downward) for $\theta>\theta^{*}$ and positive (upward) for $\theta<\theta^{*}$. Using Eq. (42), we find that the value of the radial velocity of toroidal field transport in the convective zone $v_{dens}^{light}$ (see Eqs. (40)-(42)) near low latitudes (e.g. $\theta^{*}=\arccos(0.985)\cong 10^{\circ}$; see also Fig. 7b) is almost identical to the value of the velocity (43), which was previously calculated using the toroidal field $B\approx 2\cdot 10^{5}~G$ at $0.8R_{Sun}$ (see Eq. (9) and Fig. 8b,c) in the $\Re_{dens}(\beta,\varphi)$ (see Eq. (3.13) in [93]).
@@
Thirdly, in addition to the radial profile of turbulent diffusion $\eta_{T}\approx 10^{11}~cm^{2}/s$, which is related to the diamagnetic pumping velocity $v_{dia}\approx 1.7\cdot 10^{2}~cm/s$ and the distance $2.3\cdot 10^{4}~km$ between $0.8R_{Sun}$ and near $0.85R_{Sun}$ (see the “green” line of the turbulent-toroidal magnetic field in Fig. 7b and Fig. 8b,c), we need to calculate the magnetic buoyancy velocity $v_{B}$, which is associated with a super-uniformly distributed field strength $B\sim 2\cdot 10^{5}~G$ (see Fig. 7b and Fig. 8b,c). It should be remembered that for buoyant flux tubes with super-equipartition field strength $B\sim 2\cdot 10^{5}~G\geqslant(H_{p}/a)^{1/2}B_{eq}\approx 3.16B_{eq}$ with $a\approx 0.1H_{p}$, where, as Yuhong Fan wrote (see Section 7 in [70]), $B_{eq}\approx 6.1\cdot 10^{4}~G$ is in equipartition with the kinetic energy density of the convective downflows, the magnetic buoyancy of the tubes dominates the hydrodynamic force from the convective downflows and the flux tubes can rise unaffected by convection”, but significantly affects the turbulent diamagnetic pumping!